Lab Notebook - Patrick

From OpenWetWare

Jump to: navigation, search

Contents

July

Tuesday, July 28

-Ran Analytic digests of:

  • Bdr011-from the bad plates that me and joey plated last thursday
  • all other parts from the transformation elicia did on friday.

The ccdB part might be broken, going to order oligos to do a quick change on that part tomorrow. Going to have to PCR k112122 because the sequence data came back really bad.

June

Tuesday, June 30

Today the 24 well block containing the cultures from the colonies picked from yesterdays round of colony pcr flipped over in the shaker. Goodbye better half of monday.
Analyzed the sequence data that came in today. 5 more perfects and a lot of weird insertions and deletions. Picked new colonies from the plates to grow up over night. Will be miniprepped and mapped tomorrow. Helped susan and gabi religate and transform the pcr products that were started over on account of no bands being seen. PCR products were plated.

Monday, June 29

Have missed two days of updating. Days where spent picking colonies from the bad DH10B transformations, miniprepping, running analytic gels, and sending off data to be sequenced. Saw five correct bands out of all the PCA products.

Today I analyzed the returned sequence data for parts ig302, ig306, and ig322. ig306 is perfect and the other two are most likely unusable.
Helped Susan and Gabi with colony PCR of the constructs. Picked three colonies for each construct, pcr amplified, ran on analytic gels. Saw good bands. Innoculated the overnights from good colonies in a 24 block.

Wednesday, June 24

Ran an analytic gel on the 4 PCA products: 213 239 249 and 261. 249 and 261 showed the correct bands while 213 and 239 showed very weak bands.
Cleaned up and eluted the products that showed weak bands into only 8 uL of water to attempt to concentrate the DNA. Accidently eluted into ~2 uL of ethanol.
Was forced to clean up again. Used only 30 uL of ADB and 200 uL of PE.

Did an EcoRI/BamHI digest of 213, 239, 249, and 261. Gel Purified and saw all four bands although the band for 213 was very faint.
Ligated 213, 239, 249, and 261 and all 12 of the first round of PCA products whose transformations previously failed. Also ligated a positive control 1786 EBA. Ligations on the bench at 3:17.
Heat shocked and transformed all 17 of these guys. Tubes in the 37 shaker at 4:40. Spec plates warming up.


Tuesday, June 23

Transferred oligos from the silly block it shipped in to the normal 96 well plate.
Used 12 channel, added 50 uL of water to each well. Shook. Spun Down for 2 minutes at 3000 RPM.
Transferred liquid to a normal 96 well plate. Some issues with pipetting.
Used robot to set up PCA oligo pools.

Preformed an Assembly reaction and an Amplification reaction for products:

  • 213
  • 239
  • 249
  • 261


Followed standard protocols. External Oligos were diluted from 100uM to 10uM by adding one uL of each external oligo to 9 uL water.
Reactions were cleaned up and cleaned up products added to freezer boxes. Eluted with 50 uL.

Monday, June 22

Analytic gels run by sherine, gabi, and susan showed product for all 10 PCAs ran on friday as well as for the two PCRs.
Digested all 12 products using the following protocol:

 8uL of eluted PCR product
 1uL of NEB Buffer 2
 0.5uL EcoRI
 0.5uL BamHI


Digested vector using the following protocol:

 3uL of vector DNA
 3uL of NEB Buffer 2
 1.5uL EcoRI
 1.5uL BamHI
 21ul of ddH2O


These were all digested for an hour at 37 degrees celcius.
Digests were gel purified with the following products in each well:
GEL 1:

  1. . ig1
  2. . ig27
  3. . ig57
  4. . ig79
  5. . ig113
  6. . ig121
  7. . vector
  8. . ladder


GEL 2:

  1. . vector
  2. . vector
  3. . ig129
  4. . ig137
  5. . ig145
  6. . ig171
  7. . ig201
  8. . ig207
  9. . ladder

Gel purified using the following protocol:

#cut out bands minimizing extra gel matter.
#put in ependorf tube and add 600uL of Zymo ADB buffer (brown bottle).
#heat at 55, shake and/or vortex until the gel has dissolved.
#'''If the DNA is <300bp''' add 250uL of isopropanol
#transfer into the Zymo column inside a collection tube (small clear guys)
#spin through, discard waste.
#Add 200 uL of PE buffer (which is basically 70% ethanol)
#spin through, discard waste.
#Add 200 uL of PE buffer
#spin through, discard waste.
#spin for 90 seconds, full speed to dry.
#elute with 8.5 uL of water into a fresh Eppendorf tube


Final Elution volume was 8uL for parts and 3uL for each vector.

Ligated The digestions into vector pbCA9532. Ligations on bench at 5:07. Used a "master mix" of reagents.

Friday, June 19

Did a regular zymo cleanup on the initial PCA assembly of ig1 ({<azo1653 AtD>}) and ig27 ({<Nmul_A1566 AtD>}).

   1. Add 180 uL of Zymo ADB buffer (brown bottle) to a 33uL or 50uL reaction.
   2. Transfer into the Zymo column (small clear guys)
   3. spin through, discard waste.
   4. Add 200 uL of PE or Zymo Wash buffer (which is basically 70% ethanol)
   5. spin through, discard waste.
   6. Add 200 uL of PE or Zymo Wash buffer
   7. spin through, discard waste.
   8. spin for 90 seconds, full speed to dry.
   9. elute with water into a fresh Eppendorf tube, use the same volume of water as the volume of the original reaction 


Amplified PCA assembly of ig1 and ig27 using external oligos. 1 uL oligo was added to 9 ul water to dilute oligo concentration from 100mM to 10mM. PCR program was added to thermocycler in the iGEM folder as PCA2.

Recipe

   1. 1 ul each outer oligo (10 uM)
   2. .5 ul phusion
   3. 10 ul 5x phusion buffer
   4. 5 ul 2mM dNTPs
   5. 32.5 ul H2O 
   6. 1 ul cleaned up assembly

Program

   1. 2 min initial denature at 94oC
   2. 30 sec denature at 94oC
   3. 30 sec anneal at 60oC [This should be high, as your outer oligos now have a huge overlap with the correct product]
   4. 30 sec extension at 68oC
   5. repeat 2-4 30 times total 


Labeled the cleaned up assembly as ig1 6-19 CU and ig27 6-19 CU and stored it in the yellow box in the freezer. Additionally stored the oligo dilutions as well as the pca oligo pool for ig27 and ig1

I cleaned up the Amplification reaction using the zymo cleanup protocol. The cleaned up amplification was put in the yellow box and labeled as: ig1 6-19 CU AMP and ig27 6-19 CU AMP.

Personal tools